OpenStudy (anonymous):

what are four consecutive numbers whose sum is 114

OpenStudy (anonymous):

\[\frac{114}{4}=\frac{57}{2}=28.5 \]\[27+28+29+30=114 \]

OpenStudy (triciaal):

consecutive means in order let the first number be x then the numbers are x, x+1, x+2, x+3 given the sum is 114 then add 4x + 6 = 114 x = (114-6) /4 = 27 x + 1 = 28 x + 2 =29 x + 3 = 30